OpenStudy (anonymous):

How do I simplify the derivative of ln(12x+1)

OpenStudy (anonymous):

I know that the derivative of ln(x) = 1/x I hope that helps

OpenStudy (zale101):

1/(12x+1) * d/dx(12x+1)=12*1/(12x+1)=12/(12x+1)

OpenStudy (anonymous):

Derivative of log junk times the derivative of the junk

OpenStudy (zale101):

lol

OpenStudy (anonymous):

Step 1: Define junk

OpenStudy (anonymous):

12x+1 = JUNK

OpenStudy (anonymous):

Derivative of Ln JUNK = 1/JUNK ie 1/12x+1

OpenStudy (anonymous):

Derivative of JUNK is 12 so

OpenStudy (anonymous):

Derivative of log JUNK times the derivative of the JUNK ie 12 * 1/(12x+1) =12/(12x+1)

OpenStudy (anonymous):

:)